# DrugAgent: Explainable Drug Repurposing Agent with Large Language Model-based Reasoning

### v1

@@ Abstract
Drug repurposing offers a promising avenue for accelerating drug development by identifying new therapeutic potentials of existing drugs. In this paper, we propose a multi-agent framework to enhance the drug repurposing process using state-of-the-art machine learning techniques and knowledge integration. Our framework comprises several specialized agents: an AI Agent trains robust drug-target interaction (DTI) models; a Knowledge Graph Agent utilizes the drug-gene interaction database (DGIdb), DrugBank, Comparative Toxicogenomics Database (CTD), and Search Tool for Interactions of Chemicals (STITCH) to systematically extract DTIs; and a Search Agent interacts with biomedical literature to annotate and verify computational predictions. By integrating outputs from these agents, our system effectively harnesses diverse data sources, including external databases, to propose viable repurposing candidates. Preliminary results demonstrate the potential of our approach in not only predicting drug-disease interactions but also in reducing the time and cost associated with traditional drug discovery methods. This paper highlights the scalability of multi-agent systems in biomedical research and their role in driving innovation in drug repurposing. Our approach not only outperforms existing methods in predicting drug repurposing potential but also provides interpretable results, paving the way for more efficient and cost-effective drug discovery processes. Code is available https://anonymous.4open.science/r/DrugRepurposeLLM-E0B5/.
@@ 1 Introduction
Drug repurposing, also known as drug repositioning, is a strategy for identifying new uses outside the scope of the original medical indication of existing drugs. This approach has gained significant attention due to its potential to reduce the time, cost, and risk associated with drug development (Pushpakom et al., 2019; Chen et al., 2024b; Fu et al., 2022b). With the growing complexity of biological data and the increasing availability of diverse biomedical information sources (Chen et al., 2024a; Wu et al., 2022b), there is a pressing need for innovative computational strategies that can efficiently integrate and analyze these vast datasets (Huang et al., 2021; Lu, 2018).
@@
Our framework includes three primary agents: the AI agent, the Knowledge Graph Agent, and the Search Agent. The AI agent employs the DeepPurpose package (Huang et al., 2020) to develop and refine drug-target interaction (DTI) models. The knowledge graph Agent utilizes the drug-gene interaction database (DGIdb) (Cannon et al., 2024), DrugBank (Knox et al., 2024), Comparative Toxicogenomics Database (CTD) (Davis et al., 2023), and Search Tool for Interactions of Chemicals (STITCH) (Kuhn et al., 2007) to extract and synthesize information on DTIs. The search Agent engages with biomedical literature, specifically leveraging large language models for automated data labeling and validation, a process typically resource-intensive and costly.
@@ 2 Related Works
The concept of drug repurposing has evolved significantly with the advancements in computational tools, leading to a growing body of literature that explores various methodologies. Here, we highlight key developments in the field that align with our multi-agent system approach.
Machine Learning in Drug Repurposing. Machine learning techniques have increasingly been applied to drug repurposing, demonstrating significant potential in predicting drug-disease interactions (Issa et al., 2021). Similarly, our AI Agent leverages the DeepPurpose toolkit, which has been extensively validated for its efficiency in drug-target interaction (DTI) predictions (Huang et al., 2020).
@@
Literature Search for Drug Repurposing. The automation of literature review and data extraction using AI tools, particularly large language models (LLMs), has become an essential component of modern drug discovery (Chakraborty et al., 2023). This trend is especially relevant in the field of drug repurposing, where efficient processing of vast amounts of scientific literature is crucial. Recent studies have demonstrated that LLM-based search tools can significantly enhance the efficiency and complexity of queries compared to traditional search engines (Spatharioti et al., 2023). In our framework, we leverage this approach by implementing a Search Agent that utilizes search engines as a data source for the LLM, aligning with current trends in AI-assisted literature review for drug discovery and repurposing.
Multi-Agent Systems in Biomedical Applications. While individual AI applications have shown promise, the integration of these technologies through a multi-agent system is less explored in the field of drug repurposing. However, similar multi-agent frameworks have been successfully implemented in other areas of biomedical research, such as clinical trials (Yue & Fu, 2024). These studies provide a foundation for the application of multi-agent systems in drug repurposing, suggesting that such frameworks can enhance the predictive accuracy and efficiency of computational drug discovery.
@@ 3.1 Overview of DrugAgent
Following the reasoning process, the system is capable of taking actions such as calculating repurposing scores, performing literature-based analyses, and querying knowledge graphs for novel drug-disease associations. By integrating this information through a weighted average approach, the system effectively simulates a highly knowledgeable drug repurposing researcher.
@@ 3.5 Search Agent: Information Extraction from Biomedical Literature
This simplified approach allows for rapid information gathering from publicly available sources. However, it is important to note that this method relies heavily on the quality and relevance of Google search results, and does not directly access or analyze full scientific papers or curated databases. As such, it serves as a preliminary screening tool rather than a comprehensive literature review system.
@@ 3.6 Calling External Tools
  • Comparative Toxicogenomics Database (CTD): The CTD (Davis et al., 2023) is a robust, manually curated database that provides information about chemical–gene/protein interactions, chemical–disease, and gene-disease relationships. It is particularly valuable for understanding how environmental exposures affect human health, integrating data from various species and linking chemicals, genes, diseases, phenotypes, and pathways (Chang et al., 2019; Wu et al., 2022a).
  • Search Tool for Interactions of Chemicals (STITCH): STITCH (Kuhn et al., 2007) is a database of known and predicted interactions between chemicals and proteins. It integrates information from various sources, including experimental data, predictive methods, and text-mining of scientific literature. STITCH is particularly useful for exploring the complex network of interactions between drugs, other chemicals, and proteins.
@@ 4 Experiment
4.1 Dataset
Our regression task involved predicting these Kd values given the target protein’s amino acid sequence and the drug compound’s SMILES string (a line notation for encoding molecular structures and specific instances) (Weininger, 1988). To evaluate our model’s performance, we employed three metrics: Mean Squared Error (MSE), R2 Score, and Correlation. These metrics collectively assess the accuracy of our predictions and the model’s capacity to capture the underlying patterns in DTIs.
4.2 Procedure
4.3 Implementation Details
@@
”””Specialized AI Agent for calculating DTI scores using machine learning models. Use the get_ml_score function to obtain the DTI score. Output the score in the following format: { ”ml_dti_score”: 1.0,}”””
@@
To improve the model’s prediction capabilities, we incorporate a weighted integration method within the Coordinator Agent. This method, known as score merging, aids in synthesizing the outputs from different agents into a comprehensive DTI prediction. The integration is performed using predefined weights ($\alpha$ and $\beta$) to balance the contributions of each prediction method:
$merged\_dti\_score=\alpha*ai\_score+\beta*search\_score+(1-\alpha-\beta)*kg\_score$
Here, ai_score, search_score, and kg_score represent the DTI scores from the AI Agent, Search Agent, and KG Agent, respectively. The weights $\alpha$ and $\beta$ are adjustable hyperparameters that determine the relative importance of each score in the final prediction. This approach enhances the accuracy of the model’s outputs and its ability to leverage diverse prediction methods for more robust drug repurposing predictions.

### v2

@@ Abstract
Drug repurposing, the identification of new therapeutic applications for existing drugs, offers a promising approach to accelerate drug development. This paper introduces a novel multi-agent framework that enhances the drug repurposing process by leveraging state-of-the-art machine learning techniques and integrating diverse knowledge sources. Our framework consists of three specialized agents: an AI Agent that develops robust drug-target interaction (DTI) models; a Knowledge Graph Agent that extracts DTIs from established databases such as DGIdb, DrugBank, CTD, and STITCH; and a Search Agent that engages with biomedical literature to annotate and verify computational predictions. By synthesizing outputs from these agents, our system harnesses a wide array of data sources to propose viable repurposing candidates. Preliminary results demonstrate the framework’s potential to not only predict drug-disease interactions but also to reduce the time and costs associated with traditional drug discovery methods. Our approach outperforms existing methods in predicting drug repurposing potential while providing interpretable results. This work highlights the scalability of multi-agent systems in biomedical research and their capacity to drive innovation in drug repurposing. The framework presented in this paper offers a promising pathway towards more efficient and cost-effective drug discovery processes, paving the way for accelerated therapeutic developments. Code is available https://anonymous.4open.science/r/DrugRepurposeLLM-E0B5/.
@@ 1 Introduction
Drug repurposing, also known as drug repositioning, is a strategy for identifying new uses outside the scope of the original medical indication of existing drugs. This approach has gained significant attention due to its potential to reduce the time, cost, and risk associated with drug development (Pushpakom et al., 2019; Chen et al., 2024b; Fu et al., 2022b). With the growing complexity of biological data and the increasing availability of diverse biomedical information sources (Chen et al., 2024a; Wu et al., 2022b), there is a pressing need for innovative computational strategies that can integrate and analyze these vast datasets (Huang et al., 2021; Lu, 2018).
@@
Our framework includes three primary agents: the AI agent, the Knowledge Graph (KG) Agent, and the Search Agent. The AI agent employs the DeepPurpose package (Huang et al., 2020) to develop and refine drug-target interaction (DTI) models. The knowledge graph Agent utilizes the drug-gene interaction database (DGIdb) (Cannon et al., 2024), DrugBank (Knox et al., 2024), Comparative Toxicogenomics Database (CTD) (Davis et al., 2023), and Search Tool for Interactions of Chemicals (STITCH) (Kuhn et al., 2007) to extract and synthesize information on DTIs. The search Agent engages with biomedical literature, leveraging large language models for automated data labeling and validation, a process resource-intensive and costly.
@@ 2 Related Works
The concept of drug repurposing has evolved with the advancements in computational tools, leading to a growing body of literature that explores various methodologies. Here, we highlight key developments in the field that align with our multi-agent system approach.
Machine Learning in Drug Repurposing. Machine learning techniques have been applied to drug repurposing, demonstrating significant potential in predicting drug-disease interactions (Issa et al., 2021). Similarly, our AI Agent leverages the DeepPurpose toolkit, which has been extensively validated for its efficiency in drug-target interaction (DTI) predictions (Huang et al., 2020).
@@
Literature Search for Drug Repurposing. The automation of literature review and data extraction using AI tools, particularly large language models (LLMs), has become an essential component of modern drug discovery (Chakraborty et al., 2023). This trend is relevant in the field of drug repurposing, where efficient processing of vast amounts of scientific literature is crucial. Recent studies have demonstrated that LLM-based search tools can enhance the efficiency and complexity of queries compared to traditional search engines (Spatharioti et al., 2023). In our framework, we leverage this approach by implementing a Search Agent that utilizes search engines as a data source, aligning with current trends in AI-assisted literature review for drug discovery and repurposing.
Multi-Agent Systems in Biomedical Applications. While individual AI applications have shown promise, the integration of these technologies through a multi-agent system is less explored in the field of drug repurposing. However, similar multi-agent frameworks have been implemented in other areas of biomedical research, such as clinical trials (Yue & Fu, 2024). These studies provide a foundation for the application of multi-agent systems in drug repurposing, suggesting that such frameworks can enhance the predictive accuracy and efficiency of computational drug discovery.
@@ 3.1 Overview of DrugAgent
Following the reasoning process, the system is capable of taking actions such as calculating repurposing scores, performing literature-based analyses, and querying knowledge graphs for novel drug-disease associations. By integrating this information through a weighted average approach, the system simulates a highly knowledgeable drug repurposing researcher.
@@ 3.5 Search Agent: Information Extraction from Biomedical Literature
This simplified approach allows for rapid information gathering from publicly available sources. However, it is important to note that this method relies on the quality and relevance of Google search results, and does not directly access or analyze full scientific papers or curated databases. As such, it serves as a preliminary screening tool rather than a comprehensive literature review system.
@@ 3.6 Calling External Tools
  • Comparative Toxicogenomics Database (CTD): The CTD (Davis et al., 2023) is a robust, manually curated database that provides information about chemical–gene/protein interactions, chemical–disease, and gene-disease relationships. It is valuable for understanding how environmental exposures affect human health, integrating data from various species and linking chemicals, genes, diseases, phenotypes, and pathways (Chang et al., 2019; Wu et al., 2022a).
  • Search Tool for Interactions of Chemicals (STITCH): STITCH (Kuhn et al., 2007) is a database of known and predicted interactions between chemicals and proteins. It integrates information from various sources, including experimental data, predictive methods, and text-mining of scientific literature. STITCH is useful for exploring the complex network of interactions between drugs, other chemicals, and proteins.
@@ 4 Experiment
In this section, we demonstrate the experimental results and case studies. Due to the page limit, the experimental setups, including dataset description, evaluation metrics, and implementation details, are elaborated in the Appendix.
4.1 Quantitative Results: Performance Comparison of DrugAgent and GPT-4
To evaluate the performance of DrugAgent against GPT-4 in predicting PKD scores from the BindingDB database, we compared their predictions using several key statistical metrics. Table 1 summarizes the results for 10 diverse drug-target combinations not used in parameter tuning.
Table 1: Performance comparison between DrugAgent and GPT-4 for different datasets. We report the average results of 5 independent runs and the corresponding standard deviations (in brackets). For each metric, we highlight the best method in bold. We marked the metrics where DrugAgent is significantly better than GPT-4 (pass the t-test, i.e., p-value$<$0.05) using “*”.
Metric DrugAgent GPT-4 w.o. AI Agent w.o. KG Agent w.o. Search Agent
[…]
Runtime ($\downarrow$) $\approx$5.000s $\approx$0.297s - - -
# OpenAI API tokens ($\downarrow$) $\approx$2000-3000 $\approx$100 - - -
cost of tokens ($\downarrow$) $\approx$\$0.006-\$0.027 $\approx$\$0.0014-\$0.0020 - - -
DrugAgent outperformed GPT-4 across all examined metrics. Regarding prediction accuracy, DrugAgent achieved a Mean Squared Error (MSE) of 1.836, significantly lower than GPT-4’s 13.420, indicating superior overall predictive power. The Mean Absolute Percentage Error (MAPE) and Mean Absolute Error (MAE) metrics further confirm this superiority, with DrugAgent achieving values of 0.134 and 1.081, respectively, compared to GPT-4’s 0.320 and 3.350.
DrugAgent demonstrated stronger explanatory power, evidenced by its positive R-squared (R2) value of 0.431, indicating a moderate fit to the data. In contrast, GPT-4’s R2 value of -0.460 suggests a poor fit. The Explained Variance metric reinforces this trend, with DrugAgent achieving a positive value of 0.577, while GPT-4 showed a value of -0.460. These results highlight DrugAgent’s superior ability to capture and explain the variance in the PKD score data.
DrugAgent also excels in prediction consistency. Its Max Error of 2.809 is less than half of GPT-4’s 6.490, indicating more reliable predictions across the dataset. Moreover, the strong positive correlation (0.761) between DrugAgent’s predicted and actual values, compared to GPT-4’s weak positive correlation (0.110), underscores DrugAgent’s effectiveness in capturing the underlying relationships in the data.
These results collectively showcase DrugAgent’s significantly improved predictive capabilities for BindingDB PKD scores compared to GPT-4, which is crucial for accurate binding affinity predictions in drug discovery and molecular interaction studies.
Our ablation study provides valuable insights into the importance of each component in the DrugAgent architecture. Removing the AI Agent resulted in the most severe performance degradation across all metrics, with MSE increasing dramatically to 52.349 and R2 declining to -15.228. This underscores the AI Agent’s critical role in understanding complex patterns in molecular structures and their relationship to binding affinities. The KG Agent also proved essential, as its removal led to significant performance drops, though less severe than the AI Agent. This indicates the KG Agent’s crucial contribution of domain knowledge about chemical structures and known interactions. While the Search Agent had a less dramatic impact on performance, it still contributed to the overall accuracy of the model, particularly in maintaining high correlation and explained variance.
DrugAgent achieves its superior performance with a runtime of 5.000s, which is efficient considering the complexity of its multi-agent architecture. In comparison, GPT-4’s runtime of 0.297s is faster, but at the cost of significantly reduced accuracy.
We also compared the number of OpenAI API tokens used and the associated costs. DrugAgent uses between 2000-3000 tokens per prediction, with an approximate cost of \$0.006-\$0.027, while GPT-4 uses around 100 tokens, costing $0.0014-$0.0020 per prediction. While DrugAgent has higher token usage and cost, its superior performance justifies this increased resource utilization for applications requiring high accuracy.
This comprehensive analysis highlights the synergistic interaction between the AI, KG, and Search Agents in our model architecture for accurate PKD score prediction. The results validate our design choices and provide valuable insights for future improvements and optimizations of DrugAgent. They particularly emphasize the potential for enhancing the integration and capabilities of each component to achieve even more accurate PKD score predictions in the context of the BindingDB database, while considering the balance between performance, computational resources, and cost.
4.2 Case Study
This study presents three case studies analyzing the drug repurposing potential of Topotecan for different proteins: TOP1, SLFN11, and SLC26A4. These cases represent a spectrum from known strong interactions to potentially novel connections, allowing for a comprehensive evaluation of our multi-agent system’s capabilities. In each case, the Coordinator Agent decomposed the complex problem into three subproblems, each handled by a specialized agent. The AI Agent predicted interaction potential based on structural properties, the Search Agent analyzed evidence from scientific literature, and the Knowledge Graph Agent evaluated connections in existing databases.
Case 1 examined the interaction between Topotecan and TOP1, a known strong interaction. The system calculated a final score of 11.51, confirming the established relationship. The high AI Agent score (7.65) and Knowledge Graph Agent score (1.0) aligned with the known mechanism of Topotecan as a TOP1 inhibitor, while the relatively low Search Agent score (0.27) likely reflected the well-established nature of this interaction not requiring extensive new studies.
Case 2 investigated the less understood but potentially relevant interaction between Topotecan and SLFN11. The system yielded a final score of 10.30, suggesting a noteworthy relationship. The high AI Agent score (7.36) indicated structural compatibility, while the moderate Search Agent (0.33) and Knowledge Graph Agent (0.72) scores reflected some existing evidence and established connections, albeit less comprehensive than in Case 1.
Case 3 explored an unlikely interaction between Topotecan and SLC26A4, resulting in a final score of 9.92. Despite a high AI Agent score (7.61) suggesting structural compatibility, the very low Search Agent score (0.00) indicated minimal literature evidence. The moderate Knowledge Graph Agent score (0.72) suggested some indirect connections, highlighting the system’s ability to detect potential novel interactions while also recognizing the need for careful interpretation.
These case studies demonstrated the multi-agent system’s ability to handle diverse scenarios, from known interactions to potential discoveries. The system effectively integrated various data sources and analytical methods, providing interpretable results with detailed reasoning processes. Importantly, it showed the capacity to recognize its own limitations, particularly when predictions conflicted with existing evidence.
The system offered practical insights for drug repurposing research across different levels of prior knowledge. In Case 1, it confirmed a well-established interaction. In Case 2, it suggested a potentially relevant interaction that warrants further investigation. In Case 3, it identified a possible novel interaction while appropriately flagging the low literature evidence, thus highlighting an area requiring careful experimental validation.
This comprehensive approach not only validates known interactions but also suggests potential new avenues for research. By providing nuanced analyses that consider multiple factors, the system demonstrates its value in accelerating and refining the drug repurposing process. Future work could focus on experimental validation of the system’s predictions, particularly for cases like Topotecan-SLFN11 where moderate evidence suggests a promising avenue for investigation.
5 Discussion
Our study presents a novel multi-agent system for drug repurposing that integrates machine learning, knowledge graphs, and literature search. This approach offers more robust predictions by leveraging diverse data sources and analytical methods. The system’s strength lies in its collaborative approach, which combines each agent’s specialized capabilities to evaluate complex DTIs comprehensively. The weighted integration method allows for flexible adjustment of different prediction methods, enhancing overall accuracy. However, limitations exist. The system still relies on human expertise for initial setup, limiting its scalability. It also lacks autonomous knowledge updating capabilities to keep pace with rapidly evolving pharmacological research. Furthermore, the current model does not adequately account for individual patient characteristics or drug combination effects.
Future research should focus on integrating autonomous knowledge updating mechanisms to keep the system current with the latest pharmacological research. It should also enhance the system’s ability to predict drug efficacy and handle complex drug combinations. Expanding the system’s applicability to a broader range of pharmaceutical tasks by incorporating existing models is crucial. This includes adapting the framework for Drug Response Prediction using models like drGAT (Inoue et al., 2024b), extending to DTI prediction tasks, leveraging the system’s existing knowledge graph and AI capabilities, and developing a more flexible prompting system to allow easy adaptation to various pharmaceutical tasks without major architectural changes, such as drug synergy prediction (Huang et al., 2022), drug property prediction (Xu et al., 2024), adverse drug reaction prediction (Chen et al., 2024a) or drug design (Fu et al., 2021a; 2022a).
Automating the preprocessing of complex data types, such as scRNA-seq data, is another important area for improvement. This involves implementing multiple preprocessing functions including Imputation (Inoue et al., 2024a; Inoue, 2024), Quality Control (McCarthy et al., 2017; Lu et al., 2023), and batch effect correction (Li et al., 2020; Fu et al., 2024; Haghverdi et al., 2018). Developing an optimization framework to automatically select and apply the most appropriate preprocessing methods for given datasets and integrating these preprocessing capabilities seamlessly into the existing multi-agent system are essential steps.
Validating the expanded system’s predictions through experimental studies and clinical trials to ensure real-world applicability across multiple tasks and data types (Lu et al., 2024a; Fu et al., 2023) is crucial for future development.
Appendix A Implementation
A.1 Pattern Selection and Weight Optimization for Score Integration
To determine the optimal integration method for our agent scores, we explored four different mathematical patterns and employed a constrained optimization method. We utilized a dataset comprising 3,332 drug-target pairs, each containing scores from our three specialized agents (AI, Knowledge Graph, and Search) along with corresponding ground truth interaction scores (PKD values) from the BindingDB dataset. We considered the following four patterns for score integration:
$f_{1}(\alpha,\beta,\gamma,A,B,C)$ $=\alpha A+\beta B+\gamma C,$ (3)
$f_{2}(\alpha,\beta,\gamma,A,B,C)$ $=\alpha A+(\beta B\cdot\gamma C),$
$f_{3}(\alpha,\beta,\gamma,A,B,C)$ $=(\alpha A\cdot\beta B)+\gamma C,$
$f_{4}(\alpha,\beta,\gamma,A,B,C)$ $=(\alpha A\cdot\gamma C)+\beta B,$
where $A$, $B$, and $C$ represent the scores from the AI, Knowledge Graph, and Search agents respectively, and $\alpha$, $\beta$, and $\gamma$ are the weights we are optimizing. For each pattern, we formulated the weight optimization as a constrained minimization problem:
$\underset{\alpha,\beta,\gamma}{\text{minimize}}$ $\|Y-f(\alpha,\beta,\gamma,A,B,C)\|_{2}$ (4)
subject to $\alpha,\beta,\gamma\geq 0,$
where $Y$ is the vector of ground truth PKD values and $i\in{1,2,3,4}$ corresponds to the pattern type. The optimization was performed using Sequential Least Squares Programming (SLSQP) with non-negative constraints for each pattern. After comparing the results, we found that the linear combination pattern $f_{1}$ yielded the best performance:
$f_{1}(\alpha,\beta,\gamma,A,B,C)=\alpha A+\beta B+\gamma C$ (5)
The initial optimized weights for this pattern were:
$\alpha$ $=1.24683589$ (6)
$\beta$ $=2.23513134$
$\gamma$ $=3.22163745\times 10^{-16}$
After obtaining these initial results, we rounded and adjusted the coefficients for practical implementation and to account for potential overfitting to our specific dataset. The final weights used in our merged DTI score calculation are:
$S_{\text{merged}}=1.2S_{\text{AI}}+2.2S_{\text{KG}}+0.5S_{\text{Search}}$ (7)
The selection of the linear combination pattern ($f_{1}$) and the subsequent weight adjustments reflect several important considerations:
The linear pattern provided the best fit to our data while maintaining simplicity and interpretability. The Knowledge Graph (KG) agent retains the highest weight, confirming its significant contribution to the final prediction. The AI agent continues to play a substantial role, with a weight slightly lower than the KG agent. While the initial optimization suggested a negligible role for the Search agent, we maintained its contribution at a non-trivial level (0.5) in the final model. This adjustment reflects our belief in the potential value of diverse information sources, even if not prominently represented in our current dataset.
It is worth noting that the sum of these weights (3.9) is intentionally not normalized to 1. This allows for a more flexible scaling of the final score, which can be beneficial in certain applications or when comparing across different datasets. This approach, combining pattern selection, data-driven weight optimization, and expert adjustment, ensures that our final predictions leverage the strengths of each agent while maintaining robustness and generalizability. The results suggest that a linear combination of structured knowledge from the KG agent, pattern recognition capabilities of the AI agent, and supplementary information from the Search agent contributes to accurate PKD value prediction, with the KG and AI agents playing particularly crucial roles.
A.2 Implementation Details
@@
”””Specialized AI Agent for calculating DTI scores using machine learning models. Use the get_AI_score function to obtain the DTI score. Output the score in the following format: { ”AI_score”: 1.0,}”””
@@
To improve the model’s prediction capabilities, we incorporate a weighted integration method within the Coordinator Agent. This method, known as score merging, aids in synthesizing the outputs from different agents into a comprehensive DTI prediction. The integration is performed using predefined weights ($\alpha$, $\beta$, and $\gamma$) to balance the contributions of each prediction method:
$merged\_dti\_score=\alpha*AI\_score+\beta*KG\_score+\gamma*search\_score$
Here, AI_score, search_score, and KG_score represent the DTI scores from the AI Agent, Search Agent, and KG Agent, respectively. The weights $\alpha$, $\beta$ $\gamma$ are adjustable hyperparameters that determine the relative importance of each score in the final prediction. This approach enhances the accuracy of the model’s outputs and its ability to leverage diverse prediction methods for more robust drug repurposing predictions. This parameter and the formula were defined systematically. (See A.1 in detail.)